# Minimax Decision Trees via Martingale Approximations

### v1

@@ 1 Introduction
Let $d\geqslant 1$ and $(\mathbf{X},Y)$ be a coupling (or a joint distribution) on $\mathbb{R}^{d+1}$, where $\mathbf{X}\in\mathbb{R}^{d}$ and $Y\in\mathbb{R}$. Throughout, vectors are denoted by bold symbols. We study the general problem of efficiently approximating $Y$ by conditioning on the value of $\mathbf{X}$ using finite partitions. If we adopt the $L^{2}$ norm as the minimization criterion, the problem can be formulated as
$\min_{m\in\mathcal{M}}\mathbb{E}[(Y-m(\mathbf{X}))^{2}],$ (1)
where $\mathcal{M}$ denotes some class of candidate functions defined on $\mathbb{R}^{d}$. In this paper, we focus on the case where $\mathcal{M}$ consists of piecewise constant functions. Generally speaking, the motivation is algorithmic: in many applications, such as decision trees (Breiman et al. (1984)), the goal is to provide an efficient approximation of $Y$ by a function of $\mathbf{X}$. For instance, if the conditioning on $\mathbf{X}$ is based on a partition of cardinality at most $K$, we have
=$\mathcal{M}=\bigg{\{}m:\mathbb{R}^{d}\to\mathbb{R}\,\Big{|}\,m(\mathbf{x})=\sum_{A\in\pi}m_{A}{\mathbbm{1}}_{\{\mathbf{x}\in A\}},\,\pi\in\mathfrak{P}(\mathbb{R}^{d}),\,|\pi|\leqslant K,\,m_{A}\in\mathbb{R}\bigg{\}},$
where $\mathfrak{P}(\mathbb{R}^{d})$ denotes the set of all partitions of $\mathbb{R}^{d}$. Clearly, the choice $m_{A}=\mathbb{E}[Y\mid\mathbf{X}\in A]$ is always optimal in (1), so it remains to optimize the partition $\pi$. However, finding a direct solution to (1) is generally computationally infeasible, so recursive optimizers are favorable. The common approach is to construct a sequence of nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ in $\mathfrak{P}(\mathbb{R}^{d})$ that generate the conditioning on $\mathbf{X}$, where the construction is recursive by sequentially optimizing a certain decision criterion, starting from $\pi_{0}=\{\mathbb{R}^{d}\}$. It does not hurt to consider binary partitions, as the general case follows analogously. We then formulate our main problem as follows.
(P-general). Construct nested binary partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ such that $\mathbb{E}[(Y-M_{k})^{2}]$ is small for each $k$, where
$M_{k}(\omega):=\mathbb{E}[Y\mid\mathbf{X}\in A],\quad\text{for }\mathbf{X}(\omega)\in A,\leavevmode\nobreak\ A\in\pi_{k}.$ (2)
Clearly, if $Y$ is not a function of $\mathbf{X}$, the quantity $\mathbb{E}[(Y-M_{k})^{2}]$ has a strictly positive lower bound $\mathbb{E}[\mathrm{Var}(Y\mid\mathbf{X})]$. Another obstruction arises from atoms of $\mathbf{X}$, which prevents an effective split of the partitions. The relation (2) and its analogues will be essential to this paper.
The main theme of this paper is providing solutions to (P-general) in various settings. Roughly speaking, our main contribution can be summarized as follows.
(Solving P-general). If $\mathbb{E}[Y\mid\mathbf{X}]$ is well-behaved, then modulo the lower bound $\mathbb{E}[\mathrm{Var}(Y\mid\mathbf{X})]$ and the atoms of $\mathbf{X}$, there exists an explicit construction of $\{\pi_{k}\}_{k\geqslant 0}$, not depending on the law of $(\mathbf{X},Y)$, such that $\mathbb{E}[(Y-M_{k})^{2}]$ decays exponentially in $k$, where $M_{k}$ is defined in (2).
An instance of (Solving P-general) is Theorem 9 below, where the construction is given by the cyclic MinimaxSplit algorithm detailed in Section 3.1.
In the rest of the Introduction, we illustrate the motivations of (P-general) and summarize our contribution via two concrete applications: partition-based martingale approximations and regression trees.
1.1 Partition-based martingale approximations
Let $U$ be a real-valued atomless random variable. A partition-based martingale approximation of $U$ is a discrete-time martingale $\{M_{k}\}_{k\geqslant 0}$ such that
$M_{k}(\omega):=\mathbb{E}[U\mid U\in A],\quad\text{for }U(\omega)\in A,\leavevmode\nobreak\ A\in\pi_{k},$ (3)
for some nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}$, where we assume that the partitions are binary and $\pi_{0}=\{\mathbb{R}\}$. The martingale property follows from the tower property of conditional expectations.
The connection to (P-general) is as follows. Suppose that $d=1$ (i.e., $\mathbf{X}$ is real-valued, and will be denoted by $X$ in the following), and the coupling $(X,Y)$ satisfies that $X$ is atomless and $Y$ is a strictly increasing function of $X$. Since in this case $\mathbb{E}[\mathrm{Var}(Y\mid X)]=0$ and $X$ is atomless, we expect that the approximation error $\mathbb{E}[(U-M_{k})^{2}]$ vanishes as $k\to\infty$. Moreover, each partition in $X$ bijectively maps to a partition in $Y$, and hence (2) reduces to (3). In other words, in the current setting, (P-general) is equivalent to the following problem:
(P-martingale). Given a real atomless random variable $U$, construct a partition-based martingale approximation $\{M_{k}\}_{k\geqslant 0}$ of $U$ such that $\mathbb{E}[(U-M_{k})^{2}]\to 0$ exponentially in $k$.
Motivated by a martingale embedding problem, Simons (1970) first introduced the Simons martingale and established the a.s. convergence $M_{k}\to U$ (and hence also in $L^{2}$), but did not analyze the convergence rate. Another recent motivation for studying the convergence rate of the Simons martingale arises from the construction of powerful e-values in hypothesis testing (Ramdas and Wang, 2024). Zhang et al. (2024) (Lemma 5.6) proved that if $U\in L^{2+\delta}$ for some $\delta>0$ and $\{M_{k}\}$ is the Simons martingale, then there exist $C>0$ and $r\in(0,1)$ such that $\mathbb{E}[(U-M_{k})^{2}]\leqslant Cr^{k}$, and $r<0.827$ is feasible if $U$ is bounded. Our Theorem 3(ii) improves the rate to $r=1/2$ and hence provides a tighter theoretical bound. We also show by example that the rate $r=1/2$ is optimal.
The terminology martingale approximation has been used extensively in the probability literature with different meanings. In Rüschendorf (1985), it refers to the best approximation of a random vector by a (single) martingale based on ideas from optimal transport. In Borovskikh and Korolyuk (1997) and Hall and Heyde (2014), it refers to techniques from martingale theory (such as inequalities and CLT rates) with various applications in statistics. Similar techniques are also used in the study of stationary ergodic sequences (Wu and Woodroofe, 2004; Zhao and Woodroofe, 2008) and Markovian walks (Grama et al., 2018). Note that in our setting, $Y$ is a sum of martingale differences, but whose variance is bounded, and hence one cannot apply the martingale CLT and its convergence rates.
1.2 Regression trees
Consider a regression problem where we have a data set $(\mathbf{X}_{i},Y_{i})_{1\leqslant i\leqslant N}$, where $\mathbf{X}_{i}\in\mathbb{R}^{d}$ represents the covariates (or features) and $Y_{i}\in\mathbb{R}$ represents the responses. We follow van der Vaart and Wellner (2013) and assume the regression model:
$Y_{i}=g_{*}(\mathbf{X}_{i})+\varepsilon_{i},$ (4)
where $g_{*}:\mathbb{R}^{d}\to\mathbb{R}$ is the true signal function, $\{\mathbf{X}_{i}\}_{1\leqslant i\leqslant N}$ are i.i.d. sampled from a certain law, and $\varepsilon_{i}$’s are i.i.d. random (Gaussian) errors with zero mean. In other words, the data set is sampled from some coupling $(\mathbf{X}_{*},Y_{*})$ where $Y_{*}\mid\mathbf{X}_{*}=g_{*}(\mathbf{X}_{*})+\varepsilon$ with $\mathbf{X}_{*}$ and $\varepsilon$ independent. The function $g_{*}$ can be estimated by minimizing the empirical $L^{2}$ risk (i.e., sum of squares; also called $L^{2}$ loss or mean-squared error (MSE) in this paper):
$\hat{m}_{N}=\operatorname*{arg\,min}_{m\in\mathcal{M}}\frac{1}{N}\sum_{i=1}^{N}\left(Y_{i}-m(\mathbf{X}_{i})\right)^{2},$ (5)
where $\mathcal{M}$ denotes the class of candidate functions for the mean $m$. The quantity (5) serves as an approximation to $\mathbb{E}[\left(Y_{*}-m(\mathbf{X}_{*})\right)^{2}]$. We usually establish the consistency of the estimator $\hat{m}_{N}$ by showing the convergence in probability:

### v2

@@ 1 Introduction
Tree-based predictors are classical non-parametric learners, from early systems such as THAID to the CART formalization (Morgan and Messenger, 1973; Breiman et al., 1984), with surveys charting both breadth and limitations (Gordon and Olshen, 1984; Loh, 2014) and recent work revisiting large-scale aspects (Klusowski and Tian, 2024). A persistent and practically important pathology of single-tree greedy splitting is the end-cut preference (ECP): standard impurity/variance criteria favor splits near node boundaries, especially under weak or no signal. ECP has long been observed empirically (Morgan and Messenger, 1973; Torgo, 2001; Su et al., 2024; Wang, 2025) and is now rigorously characterized in simple settings, where noisy directions drive splits toward extremes (Breiman et al., 1984; Cattaneo et al., 2022).
For a single deep tree, ECP can be harmful: aggressive boundary splits create highly unbalanced children, amplify estimation noise, and can stall meaningful refinement along informative coordinates. However, Ishwaran (2015) argues that the same tendency can be beneficial in deep random forests: when a candidate split falls on a noise variable, an end-cut keeps one child large, preserving downstream sample size so later splits (possibly on strong variables) can recover the signal; even along strong variables, ECP-like behavior can be adaptive in regions with little local signal. Thus, while ECP degrades a single tree’s stability, it may aid an ensemble by lowering tree correlation and enabling recovery via aggregation. At the same time, the forest setting introduces its own costs: ECP interacts with candidate-set multiplicity and variable-usage bias, skewing split frequencies and importance metrics, and further reducing model interpretability. Existing methods to mitigate ECP include the delta splitting rule (Morgan and Messenger, 1973), regularization of node size (Torgo, 2001), and the Smooth Sigmoid Surrogate (Su et al., 2024). We refer to Wang (2025) for a more comprehensive review of the ECP.
As illustrated in Figure 1, ECP arises from the greedy split search itself. At each node, the algorithm scans many thresholds and chooses the one with the largest impurity decay. Thresholds near the boundary isolate very few points; the sample mean of such a tiny child has a high variance, so even pure noise can create a large difference between child means and an inflated impurity reduction. Taking the maximum over many candidate cut-off points amplifies this selection bias—extremal thresholds are disproportionately likely to ‘win’.
Our focus is on explicitly mitigating ECP at the algorithmic level for trees, while retaining the transparency that makes single decision trees interpretable. Beyond prediction, this design goal is interpretability: balanced, nondegenerate splits yield partitions whose geometry is easier to understand and communicate than the heterogeneous, boundary-driven partitions that arise under ECP in deep trees and ensembles.
In this paper, we revisit ECP empirically and theoretically, clarifying the cases in which standard CART-style criteria drift toward extreme cut-off points and how this differs between single trees and forests (Breiman et al., 1984; Morgan and Messenger, 1973; Torgo, 2001; Su et al., 2024; Ishwaran, 2015; Wang, 2025; Cattaneo et al., 2022). Then, we introduce a Minimax splitting rule with a cyclic coordinate schedule that systematically discourages degenerate child nodes, thereby mitigating ECP while preserving interpretability. As a consequence, our splitting rule admits a faster convergence guarantee compared to standard CART — we establish exponential error decay under mild conditions and relate our guarantees to existing theory (Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Mazumder and Wang, 2024; Cattaneo et al., 2024). Conceptually, the analysis connects to partition-based martingale approximations (Simons, 1970; Zhang et al., 2024) and complements asymptotic and inferential developments for trees/forests (Klusowski and Tian, 2024; Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Mazumder and Wang, 2024; Cattaneo et al., 2024), as well as recent scalable partitioning frameworks (Liu et al., 2023; Luo and Pratola, 2023; Luo et al., 2024) and complexity-theoretic perspectives on hierarchical splits (Blanc et al., 2020; O’Donnell et al., 2005).
Figure 1: The comparison of the scikit-learn DecisionTreeRegressor, our MinimaxSplit tree, and VarianceSplit tree on a sinusoidal target function (in red) $Y=f(X)+\varepsilon=\sin(4\pi X)+\varepsilon$, $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1),$ and $\varepsilon\buildrel\mathrm{law}\over{\sim}\mathcal{N}(0,0.1)$, with maximum depth 4 (top row) and 5 (bottom row). We display 10 different model fits based on 10 different batches of training sets of size 100, and the average mean squared error (MSE) evaluated on a different training set across 10 batches.
2 Minimax decision trees for regression
The goal of this section is to construct MinimaxSplit decision tree algorithms that avoid ECP and attain exponential convergence. We first focus on the regression setting and we discuss the classification setting in Section 4. In CART-style regression, a decision tree greedily splits in order to minimize the sum of variances over the responses of the child nodes (Liu et al., 2023; Breiman et al., 1984); instead, we propose to greedily minimize the maximum variance among the child nodes. In Section 2.1, we rigorously introduce the algorithms (summarized in Figure 2). We study exponential rates for marginally atomless joint distributions (that may not arise from empirical measures $\mathbb{P}_{N}$) in Section 2.2. Section 2.3 builds upon this case, covering more generally the non-marginally atomless setting and, in particular, the empirical risk bounds. All proofs can be found in Appendix B.1. Throughout, we use the notation $d]=\{1,2,\dots,d\}$ for $d\in\mathbb{N}$. All splits we consider in this paper will be binary.
2.1 Algorithm formulation
We begin by recapping the greedy splitting regime in the VarianceSplit algorithm to construct an efficient decision tree. We need the following notion of a splittable set. A similar notion is also introduced by Definition 3.1 of Klusowski and Tian (2024).
Definition 2.1.
Consider a joint distribution $(\mathbf{X},Y)$ on $\mathbb{R}^{d+1}$. We say a set $A\subseteq\mathbb{R}^{d}$ is $(\mathbf{X},Y)$-unsplittable, or unsplittable, if any of the following occurs:
  • $\mathbb{P}(\mathbf{X}\in A)=0$;
  • $\mathbb{P}(\mathbf{X}\in A)>0$ and $Y\mid\mathbf{X}\in A$ is a constant;
Otherwise, we say that $A$ is $(\mathbf{X},Y)$-splittable, or splittable.
The greedy VarianceSplit construction (Section 8.4 of Breiman et al. (1984)) proceeds by introducing nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ with axis-aligned borders, in a way that sequentially and greedily minimizes the total risk $\mathbb{E}[(Y-M_{k})^{2}]$ at each step $k$ (or depth $k$, while splitting from the fixed partition $\pi_{k-1}$), where $M_{k}$ is taken as the conditional mean of $Y$ given the $\sigma$-algebra generated by the collection ${\mathbbm{1}}_{\{\mathbf{X}\in A\}},\,A\in\pi_{k}$. So, the partitions $\{\pi_{k}\}_{k\geqslant 0}$ of the input space $\mathbb{R}^{d}$ are nested and we require that each $\pi_{k}$ consists of sets that split each splittable element $A\in\pi_{k-1}$ into two hyper-rectangles with axis-aligned borders.
Formally, suppose that a hyper-rectangle $A=[a_{1},b_{1})\times\dots\times[a_{d},b_{d})$ belongs to the partition $\pi_{k-1}$ and is splittable (if $A$ is unsplittable, we do not split $A$ and $A\in\pi_{\ell}$ for every $\ell\geqslant k$). In particular, $M_{k-1}(\omega)=\mathbb{E}[Y\mid\mathbf{X}\in A]$ if $\mathbf{X}(\omega)\in A$. To find a split of $A$ at depth $k$, the VarianceSplit algorithm looks for a dimension $j\in[d]$ and covariate $x_{j}\in[a_{j},b_{j})$ such that the remaining risk is the smallest after splitting $A$ at $x_{j}$ in covariate $j$. In other words, one seeks for
$\begin{split}(j,\hat{x}_{j})&=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\Big{(}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j})\\ &\hskip 56.9055pt+\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\Big{)}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\Big{(}\mathbb{E}\big{[}(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}\big{]}\\ &\hskip 56.9055pt+\mathbb{E}\big{[}(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}\big{]}\Big{)}\\ &=:\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}(V_{\text{left}}(j)+V_{\text{right}}(j)),\end{split}$ (1)
where we break ties arbitrarily. [^1] Strictly speaking, the $\operatorname*{arg\,min}$ in (1) may not always be attained for general joint distributions $(\mathbf{X},Y)$, unless we assume that $\mathbf{X}$ is marginally either atomless or has a finite support, which will always be the case in our analysis (such as the atomicity condition in Theorem 6 below). However, in the general setting, if we allow splits in which atoms can be duplicated and assigned to both nodes (instead of only on the right node as in (1)), $\operatorname*{arg\,min}$ can be attained. In this paper, we will implicitly assume that $\operatorname*{arg\,min}$ is always attained unless otherwise stated. Consequently, the splittable set $A$ splits into the sets $A_{L}:=[a_{1},b_{1})\times\dots\times[a_{j},\hat{x}_{j})\times\dots\times[a_{d},b_{d})$ and $A_{R}:=[a_{1},b_{1})\times\dots\times[\hat{x}_{j},b_{j})\times\dots\times[a_{d},b_{d})$ to form its two descendants in the set $\pi_{k}$, and define
This leads to an associated process $\{M_{k}\}_{k\geqslant 0}$ to the joint distribution $(\mathbf{X},Y)$. [^2] The process $\{M_{k}\}_{k\geqslant 0}$ is a martingale, but we will not use this fact in the rest of the main text. To connect $M_{k}$ with the estimator $\hat{m}_{k}$ of the regression function at step $k$, we note that for any $\mathbf{x}\in\mathbb{R}^{d}$, there exists a unique $A_{\mathbf{x}}\in\pi_{k}$ containing $\mathbf{x}$, then $\hat{m}_{k}(\mathbf{x})=\mathbb{E}[Y\mid\mathbf{X}\in A_{\mathbf{x}}]$.
Figure 2: The comparison of the VarianceSplit, MinimaxSplit and cyclic MinimaxSplit trees. Here we use the notation: $A_{j,s}^{L}=A_{t}\cap\{x_{j}<s\}$$A_{j,s}^{R}=A_{t}\cap\{x_{j}\geqslant s\}$$\mathcal{S}_{j}(A_{t})$ are candidate thresholds (e.g., midpoints of sorted $x_{j}$ in $A_{t}$); $P_{t}(B)=\mathbb{P}(\mathbf{X}\in B\mid\mathbf{X}\in A_{t})$.
The MinimaxSplit algorithm.
We now introduce the MinimaxSplit algorithm. A common feature of the MinimaxSplit and VarianceSplit algorithms is that they both start from a nested sequence of partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ consisting of (at-most) binary splits into hyper-rectangles with axis-aligned borders. The two algorithms differ in the way the partition is constructed: in the MinimaxSplit setting, the split is chosen to minimize the maximum variance among the two descendants. Formally, for a splittable set $A=[a_{1},b_{1})\times\dots\times[a_{d},b_{d})\in\pi_{k-1}$, define the split location as
$\begin{split}(j,\hat{x}_{j})&=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}],\\ &\hskip 85.35826pt\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}]\Big{\}}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}),\\ &\hskip 85.35826pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\Big{\}}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}V_{\text{left}}(j),V_{\text{right}}(j)\Big{\}},\end{split}$ (4)
where we break ties arbitrarily. We call this the MinimaxSplit rule. After obtaining the partition $\pi_{k}$, define $M_{k}$ through (3).
The cyclic MinimaxSplit algorithm.
The cyclic MinimaxSplit algorithm is a variation of the MinimaxSplit algorithm. Instead of optimizing over all covariates $j\in[d]$ in (4), we cycle through the $d$ covariates as the tree grows. That is, for $k\geqslant 0$, let $j=j_{k}=1+(k\leavevmode\nobreak\ (\text{mod }d))$. For all splittable sets $A\in\pi_{k}$, we split $A$ in the $j$-th coordinate to form its two descendants at depth $k+1$. The split location is then defined as
where we break ties arbitrarily. Define $M_{k}$ by (3).
If $d=1$, the cyclic MinimaxSplit algorithm coincides with the MinimaxSplit algorithm. If $d>1$, compared to the MinimaxSplit algorithm, the cyclic MinimaxSplit algorithm is technically more tractable. Moreover, since all features are considered equally, the cyclic MinimaxSplit algorithm appears more robust in recovering the underlying geometry of the regression function, which we illustrate numerically in Section 5. On the other hand, the cyclic MinimaxSplit algorithm may be less effective in higher dimensions because the relevant dimensions may not be easily identified or reached in the first few splits, yet it avoids a pathological example in Section 2.4.
2.2 Risk bound
The goal of this section is to justify the exponential convergence under the cyclic MinimaxSplit rule. Assuming marginal non-atomicity, we show that up to universal constants and model mis-specification, the approximation error decays exponentially with rate $r^{k}$ for some explicit $r\in(0,1)$ that depends only on dimension $d$.
We note here a crucial difference from the statistics literature that our analysis of the MinimaxSplit rule first focuses on general joint distributions $(\mathbf{X},Y)$ on $\mathbb{R}^{d+1}$, where $\mathbf{X}\in\mathbb{R}^{d}$ and $Y\in\mathbb{R}$. [^3] This can be intuitively regarded as empirical measures with infinitely many samples. Empirical measures are a special case. The class of joint distributions we consider will satisfy the following marginally atomless property.
Definition 2.2.
Let $d\geqslant 1$ and $\mathbf{X}=(X_{1},\dots,X_{d})$ be an $\mathbb{R}^{d}$-valued random vector. We say that $\mathbf{X}$ (or its law) is marginally atomless if for any $j\in[d]$ and $x\in\mathbb{R}$, $\mathbb{P}(X_{j}=x)=0$.
The reason for considering general joint distributions is twofold. First, we provide more general probabilistic results that may be of interest in other settings. Second, in our derivation of the empirical risk bounds, it would be more natural and technically convenient to first study the marginally atomless setting. Although the empirical measure is not marginally atomless, only minor changes will be required to take care of the atomic case (to be discussed in Section 2.3 below).
Let us start with some simple observations on the MinimaxSplit rule. The following lemma is a simple consequence of the total variance formula and will be proved in Appendix B.1.
Lemma 2.1.
Suppose that $\mathbf{X}$ is marginally atomless and $j\in[d]$. Then the function
$\varphi_{L}:x_{j}\mapsto\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j})$ (6)
is non-decreasing and continuous. Similarly,
$\varphi_{R}:x_{j}\mapsto\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})$
is non-increasing and continuous.
Choose an arbitrary element $\hat{x}_{j}\in\operatorname*{arg\,min}_{x_{j}}\max\{\varphi_{L}(x_{j}),\varphi_{R}(x_{j})\}.$ Let us first assume that $\varphi_{L}(\hat{x}_{j})\leqslant\varphi_{R}(\hat{x}_{j})$. Define $x_{j}^{*}:=\min\{x_{j}:\varphi_{L}(x_{j})=\varphi_{R}(x_{j})\}$. By the monotonicity of $\varphi_{L}$ and $\varphi_{R}$, we have $\hat{x}_{j}\leqslant x_{j}^{*}$, and thus
$\varphi_{L}(\hat{x}_{j})\leqslant\varphi_{L}(x_{j}^{*})=\varphi_{R}(x_{j}^{*})\leqslant\varphi_{R}(\hat{x}_{j}).$ (7)
Using (7) and the fact that $\hat{x}_{j}\in\operatorname*{arg\,min}_{x_{j}}\max\{\varphi_{L}(x_{j}),\varphi_{R}(x_{j})\}$, we have
$\varphi_{L}(x_{j}^{*})=\varphi_{R}(x_{j}^{*})=\varphi_{R}(\hat{x}_{j}).$ (8)
Since the events $A_{L}=\left\{\mathbf{X}\in A,X_{j}<x_{j}^{*}\right\}$ and $A_{R}=\left\{\mathbf{X}\in A,X_{j}\geqslant x_{j}^{*}\right\}$ form a partition of $\{\mathbf{X}\in A\}$, we have by the total variance formula,
$\begin{split}\varphi_{L}(\hat{x}_{j})+\varphi_{R}(\hat{x}_{j})\leqslant\varphi_{L}(x_{j}^{*})+\varphi_{R}(x_{j}^{*})\leqslant\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}].\end{split}$ (9)
Combining (8) and (9), we have
$\begin{split}\max\Big{\{}\varphi_{L}(\hat{x}_{j}),\varphi_{R}(\hat{x}_{j})\Big{\}}=\varphi_{R}(\hat{x}_{j})\leqslant\frac{1}{2}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}].\end{split}$ (10)
The other case $\varphi_{L}(\hat{x}_{j})\geqslant\varphi_{R}(\hat{x}_{j})$ is similar. In other words, the remaining risk on each descendant is at most half the risk on the parent node. Inductively, we gain a uniform exponential decay of the maximum risk among the nodes at depth $k$:
$\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\mathrm{Var}(Y\mid\mathbf{X}\in A)=\max_{A\in\pi_{k}}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]\leqslant 2^{-k}\mathrm{Var}(Y).$ (11)
This is the key to controlling the “sizes of the nodes”. In the case of the VarianceSplit construction, such uniform control is not possible—counterexamples exist due to ECP, which we explain in the next example.
Example 1.
We revisit (a special case of) the location regression model analyzed in Cattaneo et al. (2022). Let $(X,Y)$ be an empirical measure of $\mathrm{Unif}(0,1)\times\mathcal{N}(0,1)$ on $0,1]\times\mathbb{R}$ with $n$ samples. [^4] XX is not marginally atomless, but the example can always be modified by smoothing the data without affecting its validity. Suppose that we perform VarianceSplit for one step. Theorem 4.1 of Cattaneo et al. (2022) implies that, for $n$ large enough, the split lies between the first and $\sqrt{n}$-th order statistics with probability exceeding $1/6$. With high probability (meaning 1o(1)1-o(1) as $n\to\infty$), the right node has remaining risk at least (1o(1))n(1-o(1))n while the total risk at the parent node is at most $(1+o(1))n$. By a union bound, we see that with an arbitrary $\varepsilon>0$, for sufficiently large $n$,
$\mathbb{P}\Big{(}\max_{A\in\pi_{1}}\mathbb{P}(X\in A)\mathrm{Var}(Y\mid X\in A)\geqslant(1-\varepsilon)\mathrm{Var}(Y)\Big{)}>\frac{1}{7}.$
This example also explains why assumptions such as the sufficient impurity decay (SID) condition are favorable to ensure risk decay when using the splitting criterion (1) (Chi et al., 2022; Mazumder and Wang, 2024).
To see further intuition of ECP, we consider the event $\{\mathbf{X}\in A\}$ and sub-events $A_{L}=\left\{\mathbf{X}\in A,X_{j}<\hat{x}_{j}\right\}$ and $A_{R}=\left\{\mathbf{X}\in A,X_{j}\geqslant\hat{x}_{j}\right\}$. Without loss of generality, we may assume $\mathbb{P}(\mathbf{X}\in A)=1$ and apply the total variance formula to this partition, yielding
$\mathbb{P}(\mathbf{X}\in A)\mathrm{Var}(Y)=\mathrm{Var}(Y)$ $=\mathbb{E}[\mathrm{Var}(Y\mid\sigma(A_{L}))]+\mathrm{Var}(\mathbb{E}[Y\mid\sigma(A_{L})])$
$\begin{split}&=\mathrm{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\mathrm{Var}(Y\mid A_{R})\mathbb{P}(A_{R})\\ &\hskip 56.9055pt+\left(\mathbb{E}[Y\mid A_{L}]-\mathbb{E}[Y\mid A_{R}]\right)^{2}(1-\mathbb{P}(A_{L}))\mathbb{P}(A_{L}).\end{split}$ (12)
From this, we see that when ECP occurs, $\mathbb{P}(A_{L})\approx 1$ or 0, the improvement is dominated by the first two terms of (12): $\mathrm{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\mathrm{Var}(Y\mid A_{R})\mathbb{P}(A_{R})=V_{\text{left}}(j)+V_{\text{right}}(j)$. Therefore, when ECP occurs, the sum of variances does not necessarily decay fast enough for the VarianceSplit.
However, for (cyclic) MinimaxSplit, $\max\{V_{\text{left}}(j),V_{\text{right}}(j)\}$ always displays exponential decay (11), hence “avoiding” the ECP phenomenon. In other words, the fast decay of the sum of variances is not easily guaranteed from the VarianceSplit construction due to ECP; but the exponential decay of the maximum variance can be assured by the MinimaxSplit construction, as we show in Theorem 2.
We consider the additive function class
$\mathcal{G}:=\{g(\mathbf{x}):=g_{1}(x_{1})+\dots+g_{d}(x_{d})\},$
where $\mathbf{x}=(x_{1},\dots,x_{d})$. For $g\in\mathcal{G}$, define $\left\lVert g\right\rVert_{\mathrm{TV}}$ as the infimum of $\sum_{i=1}^{d}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}$ over all such additive representations of $g$, where $\left\lVert\cdot\right\rVert_{\mathrm{TV}}$ denotes the total variation of a univariate function of bounded variation.
Theorem 2.
Suppose that $\mathbf{X}$ is marginally atomless. Let the associated $\{M_{k}\}_{k\geqslant 0}$ be constructed from the cyclic MinimaxSplit algorithm. Then uniformly for any $\delta>0$ and $k\geqslant 0$,
Remark 3 (The MinimaxSplit analogue).
Theorem 2 does not hold in general for $d\geqslant 2$ if we replace the cyclic MinimaxSplit algorithm by the MinimaxSplit algorithm, a phenomenon we illustrate in Section 2.4 below. Nevertheless, the MinimaxSplit algorithm often achieves desirable risk decay in high-dimensional settings in which cyclic MinimaxSplit is inefficient; see Appendix D.2.
The first general result for the convergence rate of the VarianceSplit algorithm without the variance decay assumption was obtained by Klusowski and Tian (2024). In a similar vein, Cattaneo et al. (2024) studied the convergence rates of the oblique CART and obtained a $1/k$ upper bound for MSE in the general setting (Theorem 1 therein) as well as an $r^{k}$ upper bound under further assumptions such as node sizes (Theorem 4 therein). The unconditional rate guarantee of $1/k$ appears significantly slower than the exponential decay rates under variance decay assumptions. (A continuous version of) Theorem 4.2 of Klusowski and Tian (2024) states that if the associated $\{M_{k}\}_{k\geqslant 0}$ is constructed from the VarianceSplit algorithm, then for each $k\geqslant 1$,
Let us compare the rates between (14) and (13). Consider $Y=g(\mathbf{X})$ with $g\in\mathcal{G}$ and $\varepsilon>0$. The following table summarizes the convergence rates of the VarianceSplit and MinimaxSplit algorithms. Note that Theorem 2 exhibits the curse of dimensionality, which cannot be circumvented in the current setting to the best of our knowledge (Cattaneo et al., 2022; Tan et al., 2022).